# and quantum information science

### v1

@@ 9.4 Bounding complexity via concentration
Another way to bound the circuit complexity of a state and, in particular, the depth necessary to prepare it is via correlations in the output probability distributions. Let $|\psi\rangle=U|0^{n}\rangle$ be a state on $n$ qubits prepared by a constant depth quantum circuit $U$. The state $|\psi\rangle$ can be used to define a probability distribution over the bistrings $x\in\{0,1\}^{n}$ via the Born rule $p_{x}\coloneqq|\langle x|\psi\rangle|^{2}$. We can show that the probability distribution is concentrated in the Boolean cube. To make this more precise, we can define the orthogonal projector
$\Pi_{\stackrel{{\scriptstyle>}}{{<}}k}\coloneqq\sum_{\begin{subarray}{c}x\in\{0,1\}^{n}\\ |x|\stackrel{{\scriptstyle>}}{{<}}k\end{subarray}}|x\rangle\langle x|,$ (143)
where $|x|$ denotes the Hamming weight of $x$, i.e. the number of 1s in the bitstring. Then, we can show that
@@
Choosing $m=c^{\prime}n$ for another constant $c^{\prime}>0$ makes this small. However, we can notice that the operator $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)$ is a sum over $lc^{\prime}n$-local terms. In particular, for a string of Hamming weight $<cn$, we have that each bitstring that has overlap with $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)|x\rangle$ has Hamming weight at most $(c+lc^{\prime})n$. Combining this limitation on the increase in Hamming weight with the approximation (145) yields
@@
Why does this property allow us to prove lower bounds on the depth necessary to prepare a state? Consider for an extreme example the GHZ state $|\mathrm{GHZ}\rangle=\frac{1}{\sqrt{2}}(|0^{n}\rangle+|1^{n}\rangle$. Clearly, the output probability distribution $p_{x}$ is a fair coin toss between $0^{n}$ and $1^{n}$ and
@@
for any $c>0$, which is in direct contradiction to the concentration property Eq. (144). But Eq. (144) is a direct consequence of the assumption that $|\mathrm{GHZ}\rangle$ was prepared by a constant depth circuit. We conclude that $|\mathrm{GHZ}\rangle$ requires a superconstant quantum circuit to prepare.
This strategy based on concentration in output probability distributions, was used in the proof of Freedman and Hasting’s NLTS conjecture [391] due to Anshu, Breuckmann and Nirkhe [392]. The NLTS theorem asserts that Hamiltonians exist such that any state with an energy density below some constant threshold is topologically ordered in the sense of Section 9.2. More precisely, there is a family of $k$-local Hamiltonians $H_{n}=\sum_{i=1}^{m}h_{i}$ acting on $n$ qubits with $m=O(n)$ and a $\varepsilon>0$ such that all states $|\psi\rangle$ with $\langle\psi|H|\psi\rangle/n\leq\varepsilon$ are topologically ordered. In particular, this result shows that topological order can persist at constant temperatures.
@@ 10 Epilogue
We conclude this review by summarizing the main results and open problems. The summary assumes the form of a dialogue, in the spirit of Sec. 3. Sofia, Sagredo, and Complexio have now read the rest of the review and are discussing their takeaways, as at a journal club.
Sagredo: I have been impatiently awaiting your arrival, as I need guidance to identify the major opportunities in quantum-complexity research. Also, it would be helpful if Sofia could summarize the main achievements obtained so far.
Sofia: I am happy to overview the main findings. The review explored multiple notions of complexity in quantum contexts: exact and approximate circuit complexities, Nielsen’s geometric approach, notions based on operator and state spreading, and definitions inspired by tensor networks. These notions have deepened our understanding of quantum systems’ evolution, including thermalization and the chaos-integrability boundary. Within the AdS/CFT correspondence, complexity is related to gravitational observables, which can elucidate the region behind a black hole’s horizon.
@@
Sofia: As a matter of fact, Refs. [316, 26, 27, 320] identified a quantitative relationship recently. Section 7.5.1 described this breakthrough: a double-scaled SYK model’s spread complexity is dual to the size of an ERB in two-dimensional theories of gravity. The duality, being exact, is a major achievement.
@@
Sofia: The duality is simplest in two bulk dimensions, and this simplicity facilitated the proof. Yet Ref. [27] reinterprets the matching in a language that should enable a generalization to higher dimensions. Additionally, insight into two-dimensional theories has advanced long-standing problems, famously including the black-hole-information paradox [393, 394, 395, 396]. The similar insight here will hopefully enable similar advancements.
@@
Sofia: They do to an extent. Consider an ensemble of theories, such as the SYK model, distinguished by random values of one or more parameters. The Krylov and spread complexities’ variances diagnose chaos [126, 165]. Whether Krylov complexity distinguishes chaotic from integrable QFTs remains an open question [186, 187]. Recent advances appear in Ref. [188, 189], which leverage von Neumann algebras and algebraic QFTs.
Complexio: Do other notions of complexity that we have heard about—apart from Krylov complexity—diagnose chaos?
Sofia: Yes, Sec. 5.3 illustrated with a finite-dimensional quantum system: as a quantum state evolves under a chaotic Hamiltonian, it follows a geodesic in Nielsen’s complexity geometry. This result depends on the geometry’s—that is, on the unitary manifold’s—having a negative average-sectional-curvature [39, 37, 96, 86]. References [82, 14] highlighted complexity geometries that satisfy this requirement. Conveniently, we can address those geometries analytically, since their curvatures are modest.
Complexio: So we can learn about chaos from quantum complexity. Can we similarly learn from complexity about the holographic dictionary?
@@
Sofia: I believe so. Whether de Sitter spacetime has a quantum dual remains an open question. If the answer is yes, then the AdS/CFT duality ports over to de Sitter spacetime, perhaps with some modifications. Consequently, we can translate quantum-gravity questions into conceptually simpler questions about quantum systems. For now, people are conjecturing such a translation, as well as conjecturing about duals of complexity [258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280, 281, 282, 283, 173, 284, 285, 286, 287, 288, 289, 290, 291, 292, 293, 294, 295, 296, 297, 298, 299, 300, 301, 302, 307, 303, 305, 304, 311, 308, 306, 309, 312, 397, 398]. Hopefully, exploring holographic complexity in de Sitter spacetime will indicate which properties the spacetime’s quantum dual should have, if it exists.
Such a development would unlock a further opportunity for future work: researchers expect the quantum dual to be an open quantum system [173, 174]. Therefore, we may learn about open quantum systems’ complexities through de Sitter holography. Several works have already begun unpacking open quantum systems’ complexities [399, 400, 401, 376, 402, 403, 404, 405].
@@
Sofia: You heard correctly. This skepticism led to the research topic that debuted in [345]: the holographic dictionary’s complexity. According to Sec. 7.6, translating between complexity and volume takes time unless quantum gravity allows for more computational power than a quantum computer. Computer scientists have proposed that the holographic dictionary might be highly complex. Their argument extends beyond the complexity=volume conjecture to other holographic complexity conjectures. Followup work focused on the duality between area and entanglement entropy [348]. Still, we are not certain for which entries the holographic dictionary is complex. Ramifications will include our ability to study quantum gravity through quantum computing.
Sagredo: These holographic concerns bring to mind your earlier claim that, even in the absence of any duality, quantum complexity merits studying within quantum information theory. Sofia has already pointed out some open quantum-information-theoretic problems centered on quantum complexity. Can you list more?
@@
Second, we should lower-bound the circuit complexities of states in families that we can describe efficiently. We already have a lower bound $\sim\log(n)n$ as discussed in Sec. 9.4. Such an endeavor may encounter obstacles that classical computer scientists face. More precisely, pseudorandom functions could hinder the establishment of these circuit-complexity lower bounds. The reason, roughly, is the following: imagine a quantity that distinguishes low-complexity-functions from high-complexity-functions. This quantity must, itself, be hard to compute. Otherwise, one could easily use the quantity to distinguish pseudorandom functions (low complexity) from random functions (high complexity). The resulting obstacles for lower-bounding circuit complexity, we call natural proof barriers [58]. We might quantize these natural proof barriers by strengthening a recent construction of pseudorandom unitaries [347].
Third, researchers expect that some local Hamiltonians’ low-energy states have superpolynomial circuit complexities. However, proving an unconditional superpolynomial lower bound would come close to separating the complexity classes $\mathrm{QMA}$ (quantum Merlin-Arthur) and $\mathrm{QCMA}$ (quantum-classical Merlin-Arthur) [406]. QMA contains decision problems that one can verify with a quantum proof; and QCMA, decision problems that a quantum computer can verify with a classical proof. Unconditionally separating QMA and QCMA, one would achieve a monumental breakthrough in theoretical computer science by also separating the classes P and PSPACE. Compare Ref. [407] for details. A more modest goal is to separate QMA from QCMA “relative to an oracle”. An oracle is a black box that solves a powerful problem in one step. Often, unconditional statements about complexity classes can be made if all Turing machines are given access to such an oracle. The problem of separating QMA from QCMA relative to increasingly weaker oracles was subject to significant efforts in the past years [408, 406, 409, 410, 411, 412, 413, 414].
@@ Acknowledgments
We would like to thank Brian Swingle and Beni Yoshida for their involvement in the early stages of this review. We are grateful to our collaborators on the topic of quantum complexity who significantly contributed to both the knowledge development described here, as well as to our viewpoints on open questions in this area. Furthermore, we thank Igal Arav, Javier Magan, and Tim Schuhmann for valuable discussions during the writing process. We especially thank Saskia Demulder, Damian Galante, Tim Schuhmann, Tal Schwartzman and Zixi Wei for insightful discussions and comments on the draft. The authors gratefully acknowledge the organizers and participants of the workshop “EuroStrings 2024” in Southampton, United Kingdom, where part of this work was carried out. The work of S. B. was supported by the INFN grant “Gauge Theories and Strings” (GAST), via a research grant on “Holographic dualities, quantum information and gravity.” V. B. is supported in part by the DOE through DE-SC0013528 and QuantISED grant DE-SC0020360, and in part by the Eastman Professorship at Balliol College, University of Oxford. The work of S. B. and S. C. was supported by the Israel Science Foundation (grant No. 1417/21), by the German Research Foundation through the German-Israeli Project Cooperation (DIP) grant “Holography and the Swampland,” by Carole and Marcus Weinstein through the BGU Presidential Faculty Recruitment Fund, and by the ISF Center of Excellence for theoretical high-energy physics. S. C. is also supported by the ERC starting Grant dSHologQI (project number 101117338). P. C. is supported by NCN Sonata Bis 9 2019/34/E/ST2/00123 grant and by an ERC Consolidator grant (number: 101125449/acronym: QComplexity). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. N. Y. H. received support from the National Science Foundation (QLCI grant OMA-2120757). J. H. is supported by the Harvard Quantum Initiative (HQI). M. P. H. has received support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant number: 101089093 / project acronym: High-TheQ).

### v2

@@ 9.4 Bounding complexity via concentration
Another way to bound the circuit complexity of a state and, in particular, the depth necessary to prepare it is via correlations in the output probability distributions. Let $|\psi\rangle=U|0^{n}\rangle$ be a state on $n$ qubits prepared by a constant depth quantum circuit $U$. The state $|\psi\rangle$ can be used to define a probability distribution over the bitstrings $x\in\{0,1\}^{n}$ via the Born rule $p_{x}\coloneqq|\langle x|\psi\rangle|^{2}$. We can show that the probability distribution is concentrated in the Boolean cube. To make this more precise, we can define the orthogonal projector
$\Pi_{<k}\coloneqq\sum_{\begin{subarray}{c}x\in\{0,1\}^{n}\\ |x|<k\end{subarray}}|x\rangle\langle x|,$ (157)
and $\Pi_{>k}$ defined similarly. Here $|x|$ denotes the Hamming weight of $x$, i.e. the number of 1s in the bitstring. Then, we can show that
@@
Choosing $m=c^{\prime}n$ for another constant $c^{\prime}>0$ makes this small. However, we notice that the operator $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)$ is a sum over $lc^{\prime}n$-local terms. In particular, for a string of Hamming weight $<cn$, we have that each bitstring that has overlap with $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)|x\rangle$ has Hamming weight at most $(c+lc^{\prime})n$. Combining this limitation on the increase in Hamming weight with the approximation (159) yields
@@
Why does this property allow us to prove lower bounds on the depth necessary to prepare a state? Consider for an extreme example the GHZ state $|\mathrm{GHZ}\rangle=\frac{1}{\sqrt{2}}(|0^{n}\rangle+|1^{n}\rangle)$. Clearly, the output probability distribution $p_{x}$ is a fair coin toss between $0^{n}$ and $1^{n}$ and
@@
for any $c>0$, which is in direct contradiction to the concentration property in Eq. (158). But Eq. (158) is a direct consequence of the assumption that $|\mathrm{GHZ}\rangle$ was prepared by a constant depth circuit. In fact, we can show that circuits of depth $\Omega(\log(n))$ are required to prepare $|\mathrm{GHZ}\rangle$.
This strategy based on concentration in output probability distributions,was used in the proof of Freedman and Hasting’s NLTS (“No Low-Energy Trivial States”) conjecture [462] due to Anshu, Breuckmann and Nirkhe [463]. The NLTS theorem asserts that Hamiltonians exist such that any state with an energy density below some constant threshold is topologically ordered in the sense of Section 9.2. More precisely, there is a family of $k$-local Hamiltonians $H_{n}=\sum_{i=1}^{m}h_{i}$ acting on $n$ qubits with $m=O(n)$ and an $\varepsilon>0$ such that all states $|\psi\rangle$ with $\langle\psi|H|\psi\rangle/n\leq\varepsilon$ are topologically ordered. In particular, this result shows that topological order can persist at constant temperatures.
@@ 10 Epilogue
Sofia, Sagredo, and Complexio have now read the the review and are discussing their takeaways at a journal club.
Sagredo: I have been impatiently awaiting your arrival, as I need guidance to identify the major opportunities in quantum-complexity research. Also, it would be helpful if Sofia could summarize the main achievements so far.
Sofia: I am happy to overview the main findings. The review explored several notions of quantum complexity: exact and approximate circuit complexities, Nielsen’s geometric approach, notions based on operator and state spreading, and definitions inspired by tensor networks. These notions have deepened our understanding of the evolution of quantum systems, including thermalization and differences between chaos and integrability. Within the AdS/CFT correspondence, complexity appears to be related to certain gravitational observables, which can perhaps elucidate the physics of the region behind black hole horizons.
@@
Sofia: As a matter of fact, Refs. [33, 34, 35, 370] built on [32] to identify a quantitative relationship recently. Section 7.5.1 described this breakthrough: a double-scaled SYK model’s spread complexity is dual to the size of an ERB in two-dimensional theories of gravity. The duality, being exact, is a major achievement.
@@
Sofia: The duality is simplest in two bulk dimensions, and this simplicity facilitated the proof. Yet [35] reinterprets the matching in a language that should enable a generalization to higher dimensions. Additionally, insight into two-dimensional theories has advanced long-standing problems, famously including the black-hole-information paradox [464, 465, 466, 8]. The similar insight here will hopefully enable similar advancements.
@@
Sofia: They do to an extent. Consider an ensemble of theories, such as the SYK model, distinguished by random values of one or more parameters. Variances associated to Krylov and spread complexities diagnose chaos [23, 24, 184]. Furthermore, individual chaotic theories, even if we do not consider an ensemble, have more regular tridiagonal spectra (also called Lanczos spectra) than integrable ones. In fact, there is a precise conjecture for the functional form of the tridiagonal spectral covariances in chaotic theories [24]. Whether and how precisely Krylov complexity distinguishes chaotic from integrable QFTs remains an open question [205, 206], although again the key distinction should lie in the fluctuations of the tridiagonal spectrum of the Hamiltonian. Recent advances appear also in Refs. [207, 208], which leverage von Neumann algebras and algebraic QFTs.
Complexio: Do the other notions of complexity that we have heard about—apart from Krylov complexity—diagnose chaos?
Sofia: Yes, Sec. 5.3 discussed constraints on the complexity metric for finite-dimensional quantum system so that Nielsen complexity shows distinguishing features for chaotic dynamics. Similarly, Sec. 5.4 established physical criteria that lead to long-term Nielsen complexity growth in metrics that penalize complex versus simple generators. Systems that satisfy these criteria, which include some that show spectral features of chaos like level repulsion, show linear growth of Nielsen complexity for exponential time, at least up to possible obstructions arising from the global structure of the unitary group that are hard to study. This long-term complexity growth has its roots in the negative average sectional curvature of the complexity geometry on the unitary group manifold [50, 108, 112, 101]. References [97, 19] discussed various classes of complexity geometries that penalize complex operations as desired. These geometries have a negative average sectional curvature, and are constructed to facilitate analytical computation of geodesics as desired for the study of Nielsen complexity.
Complexio: So we can learn about chaos from quantum complexity. Can we also learn from complexity about the holographic dictionary?
@@
Sofia: Perhaps. Whether de Sitter spacetime has a quantum dual remains an open question. If the answer is yes, then the methods of the AdS/CFT duality may port over to de Sitter spacetime, perhaps with some modifications. If so, we can translate quantum-gravity questions into conceptually simpler questions about non-gravitating quantum systems. For now, people are conjecturing such a translation, as well as conjecturing about duals of complexity [305, 306, 307, 308, 309, 310, 313, 314, 315, 316, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 327, 328, 329, 330, 331, 332, 192, 333, 334, 335, 336, 337, 338, 339, 278, 340, 341, 342, 343, 344, 345, 346, 347, 348, 349, 350, 358, 353, 355, 354, 362, 359, 357, 360, 364, 467, 468]. Hopefully, exploring holographic complexity in de Sitter spacetime will indicate which properties the spacetime’s quantum dual should have, if it exists.
Such a development would unlock further opportunities for future work: some researchers expect the quantum dual of de Sitter space to be an open quantum system [192, 193]. If this is the case, we may learn about complexity of dynamics in open quantum systems through de Sitter holography. Several works have already begun unpacking complexity in open quantum systems [469, 470, 471, 429, 472, 473, 474, 475, 476, 477, 478, 479].
@@
Sofia: You heard correctly. This skepticism led to the research topic that debuted in [397]: the holographic dictionary’s complexity. According to Sec. 7.7, translating between complexity and volume takes a lot of time unless quantum gravity allows for more computational power than a quantum computer. Computer scientists have proposed that the holographic dictionary might be highly complex. Their argument extends beyond the complexity=volume conjecture to other holographic complexity conjectures. Followup work focused on the duality between area and entanglement entropy [405]. Still, we are not certain for which entries the holographic dictionary is complex. Ramifications will include our ability to study quantum gravity through quantum computing.
Sagredo: These holographic concerns bring to mind your earlier claim that, even in the absence of any duality, quantum complexity merits studying within quantum information theory. Sofia, you have already pointed out some open quantum-information-theoretic problems centered on quantum complexity. Can you list more?
@@
Second, we should lower-bound the circuit complexities of states in families that we can describe efficiently. We already have a lower bound $\sim n\,\log(n)$ as discussed in Sec. 9.4. Such an endeavor may encounter obstacles that classical computer scientists face. More precisely, pseudorandom functions could hinder the establishment of such circuit-complexity lower bounds. The reason, roughly, is the following: imagine a quantity that distinguishes low-complexity-functions from high-complexity-functions. This quantity must, itself, be hard to compute. Otherwise, one could easily use the quantity to distinguish pseudorandom functions (low complexity) from random functions (high complexity). The resulting obstacles for lower-bounding classical circuit complexity are called natural proof barriers [70]. We might develop the quantum analog of such natural proof barriers by strengthening a recent construction of pseudorandom unitaries [399].
Third, researchers expect that the low-energy states of some local Hamiltonians have superpolynomial circuit complexities. However, proving an unconditional superpolynomial lower bound would come close to separating the complexity classes $\mathrm{QMA}$ (Quantum Merlin-Arthur) and $\mathrm{QCMA}$ (Quantum-Classical Merlin-Arthur) [480]. QMA contains decision problems that one can verify with a quantum proof; and QCMA, decision problems that a quantum computer can verify with a classical proof. Unconditionally separating QMA and QCMA, one would achieve a monumental breakthrough in theoretical computer science by also separating the classes P and PSPACE (see [481] for details). A more modest goal is to separate QMA from QCMA “relative to an oracle”. An oracle is a black box that solves a powerful problem in one step. Often, unconditional statements about complexity classes can be made if all Turing machines are given access to such an oracle. The problem of separating QMA from QCMA relative to increasingly weaker oracles was the subject of significant efforts in the past years [482, 480, 483, 484, 485, 486, 487, 488].
@@ Acknowledgments
We would like to thank Brian Swingle and Beni Yoshida for their involvement in the early stages of this review. We are grateful to our collaborators on the topic of quantum complexity who significantly contributed to both the knowledge development described here, as well as to our viewpoints on open questions in this area. Furthermore, we thank Igal Arav, Andrea Legramandi, Javier Magan, Tim Schuhmann and Paolo Stornati for valuable discussions during the writing process. We especially thank Saskia Demulder, Damian Galante, Tim Schuhmann, Tal Schwartzman and Zixi Wei for insightful discussions and comments on the draft. The authors gratefully acknowledge the organizers and participants of the workshop “EuroStrings 2024” in Southampton, United Kingdom, where part of this work was carried out. The work of S. B. was supported by the INFN grant “Gauge Theories and Strings” (GAST), via a research grant on “Holographic dualities, quantum information and gravity.” V. B. is supported in part by the DOE through DE-SC0013528 and QuantISED grant DE-SC0020360, and in part by the Eastman Professorship at Balliol College, University of Oxford. The work of S. B. and S. C. was supported by the Israel Science Foundation (grant No. 1417/21), by the German Research Foundation through the German-Israeli Project Cooperation (DIP) grant “Holography and the Swampland,” by Carole and Marcus Weinstein through the BGU Presidential Faculty Recruitment Fund, and by the ISF Center of Excellence for theoretical high-energy physics. S. C. is also supported by the ERC starting Grant dSHologQI (project number 101117338). P. C. is supported by NCN Sonata Bis 9 2019/34/E/ST2/00123 grant and by an ERC Consolidator grant (number: 101125449/acronym: QComplexity). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. N. Y. H. received support from the National Science Foundation (QLCI grant OMA-2120757). J. H. is supported by the Harvard Quantum Initiative (HQI). M. P. H. has received support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant number: 101089093 / project acronym: High-TheQ).